# A Closer Look into Mixture-of-Experts in Large Language Models

### v2

@@ 5.2 Norms of Expert Outputs and Gate Scores
Mixtral. We found that the two experts chosen by the gate usually output feature vectors with the highest norms, which reveals that the norm might be one of the key factors in gate decisions. This finding agrees with the router’s design in CompeteSMoE Pham et al. (2024), which selects experts based on their output norms. It also helps explain why the outputs of the chosen Mixtral and DeepSeek experts tend to be more alike (§ 5.1). In Fig. 4, we observe that the gate scores assigned to the top-1 experts are usually much higher than those of the others, including the second place. This demonstrates that the gate is learned to strengthen the confidence of its decision during the training process. On the other hand, the deeper the layer, the larger the norm, which is similar to the growth in standard models Shleifer et al. (2021).
DeepSeek. In contrast to the observation about Mixtral experts, the gate decision appears to depend less obviously on the output norms of DeepSeek experts. However, the top-1 experts often score much higher than the remaining candidates. The magnitude of the norms increases with depth, although the increment is less pronounced than in Mixtral. In the last layer, the variance of norms becomes greater.
Grok. While the scores of the top-1 experts are higher than those of the others, no correspondence between the norms and the gate scores is observed. One possible reason could be the relatively low activation ratios of GeLU (see Append H), which may lead to a weaker dependence on the norm for gate decisions. Besides, unlike Mixtral and DeepSeek, the magnitude of the norms hardly changes across depth, and some of the norm values can be less than 1, which is rare in the other two models.
@@ 5.3 Summary
The observations of dynamic behaviours are concluded below: i) The outputs of Mixtral and DeepSeek experts in deep (last) layers are less (much) alike. This can be seen in the heat maps for both the short (Fig. 2) and long (Fig. 3) inputs. ii) The average heat maps of expert outputs resemble the neuron-level similarity graphs (Fig. 3), implying that weight similarity measurements can reflect output similarity. iii) Grok experts exhibit high output similarity (Fig.2), likely due to their larger sizes. iv) For Mixtral and DeepSeek, experts generating feature vectors with larger norms tend to receive higher gate scores, as shown in Fig. 4. We further verified this observation in Fig. 10.
@@ 6 Discussion
Neuron-level experts. Intuitively, the gate embedding determines expert selection while $W_{\text{gate}}$ is responsible for choosing which neurons to activate. Meanwhile, we found that the similarities of the gate embedding and $W_{\text{gate}}$ show association. This implies that neurons may function as more fine-grained experts. Therefore, operations on experts, such as division, construction, and composition, should be further studied at the micro level. For instance, MoEfication Zhang et al. (2021) and EMoE Qiu et al. (2024a) construct MoE experts by splitting the MLP layers of a dense model, suggesting our findings from a similar perspective.
@@ 7 Related Work
Most existing works analyze MoE from the router’s perspective by observing expert selections. Early works have observed the unstable choices in the router Zuo et al. (2021); Chi et al. (2022); Dai et al. (2022). More recent studies find the standard routers do not show clear specialization at the domain level Jiang et al. (2024); Dai et al. (2024) and primarily route based on token ID instead of high-level semantics Xue et al. (2024). Shi et al. (2024) shows that Top-2 and Rank-k routing result in different model behaviours and propose a new self-contrast decoding method to determine the next-token distribution based on this finding.
@@ 8 Conclusion
In this paper, we initially attempt to investigate the inner working mechanisms of MoEs by studying the parameters and outputs of three different MoE models. We summarize our empirical observations and propose practical suggestions across various aspects. While it is premature to conclude whether MoEs genuinely learn heterogeneous experts, some of our experiments indicate that specific architectural designs (e.g., the number of experts) and training frameworks may facilitate expert specialization. We hope this work can provide inspiring insights and serve as a valuable foundation for future research on MoE and other modular architectures.
@@ Appendix A Model Selection
Our experiments are conducted on Mixtral 8x7B, DeepSeekMoE, and Grok-1. We choose these models due to their widespread use and impressive performance across various domains. Additionally, these models are complementary in several crucial attributes, such as training scheme, activation functions, top-k settings, and the number of experts, as listed in Tab 1 and Tab 4. This allows for a comparative analysis with controlled variables and encompasses a wide range of parameter sizes, from rather small (16B) to relatively huge (314B). Hence, we believe that the findings derived from these three models are fairly robust, despite the limited number of models examined.
@@ Appendix B Extended Related Work
MoE LLMs. MoEs have garnered significant attention in recent years due to their ability to efficiently scale model capacity with minimal computational overhead. Most current transformer-based MoE LLMs adopt a typical architecture design that replaces the original FFN with multiple expert networks and a sparse gating network Wei et al. (2024a); Wu et al. (2024); Dai et al. (2024); Xue et al. (2024); Jiang et al. (2024); Zoph et al. (2022). JetMoE Shen et al. (2024) and ModuleFormer Shen et al. (2023) incorporate Mixture of Attention Heads Zhang et al. (2022) into their model, achieving further sparsity. A recent survey Cai et al. (2024) provides a comprehensive review of both the algorithmic and system design aspects of MoEs. For this study, we select three representative candidates among current open-sourced MoE LLMs for analysis to gain intriguing insights.
@@ C.1 Matrix-level
Figure 6: Projection of expert matrices in 2D space. Each layer contains three graphs, corresponding to $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, respectively. For DeepSeek, the indices of the removed outliers are listed on top of each graph.
@@ C.2 Neuron-level
Figure 7: Projection of expert neurons in 2D/3D space. Each layer contains three graphs, corresponding to $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, respectively.
To project the neurons into a 2D or 3D space, each row vector of $W_{\text{up}}$ and $W_{\text{gate}}$, or each column vector of $W_{\text{down}}$, is treated as a single data point. Standardization is then applied, following by PCA. The visualization of the principal components is illustrated in Fig. 7. Different colors refer to neurons belonging to different experts.
Common. The vast majority of neurons gather in the low-dimensional space. In some layers, the distribution of neurons forms a special shape, such as a cross or a thick line, which appears the most often for $W_{\text{down}}$, followed by $W_{\text{up}}$, and finally $W_{\text{gate}}$. Compared to ellipses, these shapes indicate that the neurons are relatively more similar.
@@ Appendix D Averaging Expert Neurons
To investigate expert correlation at the neuron level, the averaging approach simply averages the rows (for $W_{\text{up}}$ and $W_{\text{gate}}$) or the columns (for $W_{\text{down}}$) of the weight matrices and then calculates the similarity of the resulting vectors across experts. Fig. 8 displays the graphs.
Common. The heat maps of $W_{\text{up}}$ and $W_{\text{down}}$ are nearly identical to those presented in § 4.1.1. Yet the similarities of $W_{\text{gate}}$ significantly increase.
Mixtral and Mistral. The dark crosses sometimes disappear. In the figures for $W_{\text{gate}}$, the similarities between the experts and the Mistral FFN are often lower than the similarities among the experts themselves (i.e., $S_{\text{ee}}>S_{\text{ef}}$), which is contrary to previous observations. This can happen if the expert neurons in different positions are alike. For instance, given three vectors $f=(0,0)$, $e_{1}=(1,0)$, and $e_{2}=(0,1)$, the vector similarity $S_{e_{1}e_{2}}$ is lower than $S_{e_{1}f}$ and $S_{e_{2}f}$. If averaging the elements, we have $\bar{f}=(0)$, $\bar{e}_{1}=(0.5)$, and $\bar{e}_{2}=(0.5)$, then $S_{e_{1}e_{2}}$ becomes the highest.
DeepSeek. The growth of the $W_{\text{gate}}$ similarity values is directly proportional to the layer depth.
Gork. In the heat map of $W_{\text{gate}}$, dark crosses frequently appear in various positions.
@@ Appendix E Gate Embedding
Layer Mixtral DeepSeek Grok
0 0.82 0.89
1 -0.44 0.75 -0.10
2 0.26 0.78 -0.28
3 0.54 0.71 0.66
4 0.48 0.77 0.52
5 0.70 0.77 0.37
6 0.84 0.69 0.28
7 0.74 0.73 0.17
8 0.42 0.66 0.51
9 0.66 0.66 0.84
10 0.53 0.63 0.28
11 0.32 0.60 0.30
12 0.14 0.54 0.46
13 0.51 0.60 0.14
14 0.66 0.56 0.00
15 0.40 0.58 0.54
16 0.39 0.53 0.32
17 0.53 0.55 0.30
18 0.35 0.57 0.10
19 0.17 0.57 -0.17
20 0.51 0.58 0.24
21 0.63 0.62 0.58
22 0.36 0.62 0.46
23 0.51 0.62 0.14
24 0.48 0.68 0.10
25 0.66 0.62 0.00
26 0.81 0.58 -0.10
27 0.63 0.46 -0.26
28 0.73 -0.66
29 0.75 -0.41
30 0.84 -0.83
31 0.57 -0.76
32 -0.24
33 -0.53
34 -0.46
35 0.14
36 0.17
37 -0.46
38 -0.17
39 -0.26
40 -0.70
41 0.17
42 0.00
43 -0.17
44 -0.22
45 0.14
46 -0.47
47 -0.44
48 -0.17
49 -0.14
50 0.17
51 0.22
52 0.10
53 0.33
54 -0.24
55 -0.57
@@
Table 5: Pearson correlation coefficients ($R$) of the paired dataset $(X,Y_{\text{gate}})$.
Since each row vector of the gate embedding determines the gate decision, there may be a relationship between the gate embedding matrix and that of the experts. To investigate this, we measure the similarities between the gate embedding vectors and compare them with the average heat maps presented in § 4.1.2. The qualitative analysis of the combined graphs shown in Fig. 8 is detailed in this section. The table containing the $R$ values for each layer (Tab. 5) is appended at the end.
Mixtral. Focusing on the heat maps of the gate embedding, the similarities typically range from 0.2 to 0.4, with a noticeable increase in the last layer. Moreover, dark crosses are rarely found. Surprisingly, the patterns in the heat maps of gate embeddings and expert neurons in $W_{\text{gate}}$ are partially alike in some layers. This implies that the way a gate selects experts might be relevant to how an expert activates its neurons.
DeepSeek. Unlike the almost all-zero heat maps of $W_{\text{up}}$ and $W_{\text{down}}$, the similarities of gate neurons sometimes exceed 0.4. In addition, the heat maps of gate embeddings and $W_{\text{gate}}$ show similar patterns. However, the overall similarities of gate embeddings decrease with depth while the similarities of $W_{\text{gate}}$ gradually grow. This indicates that as the layer depth increases, the gate “looks” at the input feature in more diverse ways when assigning scores to different experts, even as the neuron activations of the experts become more similar.
Gork. Both dark and bright crosses commonly exist in the heat maps of the gate embedding, whose patterns are similar to those of $W_{\text{gate}}$. Specially, their patterns show opposite color tendency (i.e., deep color positions in one heat map becomes light color in another) starting form the intermediate layers. The similarities of the gate embedding decrease when the layer depth increases, except for the last few layers.
@@ Appendix F Additional Datasets
Figure 9: Average similarity heat maps of expert output features using (1) the long sequence, (2) WikiText-103, (3) GSM8K, and (4) Magicoder-Evol-Instruct-110K. The tick numbers refer to the expert indices.
@@ Appendix H Intermediate States of Experts
While § 5.1 focused on the final outputs of experts, we continue our analysis here by examining their intermediate outputs to examine the inner states of the experts. Given an input $x$, the intermediate state of an expert refers to the output of $\operatorname{act}(W_{\text{gate}}x)$, which is a $d_{\text{hid}}$-dimensional vector. These intermediate vectors control the activation of neurons, so we simply record them for analysis with the short input used. Mixtral, Mistral, and DeepSeek utilize SiLU as the activation function, while Grok adopts GeLU. Fig. 12 depicts the magnitude of the vectors for Mixtral across three tokens.
@@
Mixtral and Mistral. For a single token, we found that, on average, the absolute activation value of 99.6% elements in each expert exceeds 0.001 after applying the SiLU activation function. This high ratio indicates that the vast majority of neurons in an expert are activated. In Fig. 12, some vertical lines across all experts are commonly found, meaning that the $W_{\text{gate}}$ matrices of different experts assign similar activation values to neurons with the same indices. In addition, the magnitude of the intermediate states grows along with layer depth, which aligns with the observation in § 5.1.

### v3

@@ 5.2 Norms of Expert Outputs and Gate Scores
Mixtrals. We found that the two experts chosen by the gate usually output feature vectors with the highest norms, which reveals that the norm might be one of the key factors in gate decisions. This finding agrees with the router’s design in CompeteSMoE Pham et al. (2024), which selects experts based on their output norms. It also helps explain why the outputs of the chosen Mixtrals and DeepSeek experts tend to be more alike (§ 5.1). In Fig. 4, we observe that the gate scores assigned to the top-1 experts are usually much higher than those of the others, including the second place. This demonstrates that the gate is learned to strengthen the confidence of its decision during training. On the other hand, the deeper the layer, the larger the norm, which is similar to the growth in standard models Shleifer et al. (2021).
DeepSeek. In contrast to the observation about Mixtrals’ experts, the gate decision appears to depend less obviously on the output norms of DeepSeek experts. However, the top-1 experts often score much higher than the remaining candidates. The magnitude of the norms increases with depth, although the increment is less pronounced than in Mixtrals. In the last layer, the variance of norms becomes greater.
Grok. While the scores of the top-1 experts are higher than those of the others, no correspondence between the norms and the gate scores is observed. One possible reason could be the relatively low activation ratios of GeLU (see Append H), which may lead to a weaker dependence on the norm for gate decisions. Besides, unlike Mixtrals and DeepSeek, the magnitude of the norms hardly changes across depth, and some of the norm values can be less than 1, which is rare in the other two models.
@@ 5.3 Summary
The observations of dynamic behaviours are concluded below: i) The outputs of Mixtrals and DeepSeek experts in deep (last) layers are less (much) alike. This can be seen in the heat maps for both the short (Fig. 2) and long (Fig. 3) inputs. ii) The average heat maps of expert outputs resemble the neuron-level similarity graphs (Fig. 3), implying that weight similarity measurements can reflect output similarity. iii) Grok experts exhibit high output similarity (Fig.2), likely due to their larger sizes. iv) For Mixtrals and DeepSeek, experts generating feature vectors with larger norms tend to receive higher gate scores, as shown in Fig. 4. We further verified this observation in Fig. 10.
@@ 6 Discussion
Neuron-level experts. Intuitively, the gate embedding matrix $W_{g}$ determines expert selection while $W_{\text{act}}$ is responsible for choosing which neurons to activate. Meanwhile, we find that the similarities of $W_{g}$ and of $W_{\text{act}}$ show association. This implies that neurons may function as more fine-grained experts. Therefore, operations on experts, such as division, construction, and composition, should be further studied at the micro level. For instance, MoEfication Zhang et al. (2021) and EMoE Qiu et al. (2024a) construct MoE experts by splitting the MLP layers of a dense model, suggesting our findings from a similar perspective.
@@ 7 Related Work
Most existing works analyze MoE from the router’s perspective by observing expert selections. Early works have observed the unstable choices in the router Zuo et al. (2021); Chi et al. (2022); Dai et al. (2022). More recent studies find the standard routers do not show clear specialization at the domain level Jiang et al. (2024); Dai et al. (2024) and primarily route based on token ID instead of high-level semantics Xue et al. (2024). Shi et al. (2024) shows that Top-2 and Rank-k routing result in different model behaviours and proposes a new self-contrast decoding method to determine the next-token distribution based on this finding.
@@
As mentioned in previous sections, several existing works share some relevance to our findings, and thus can be seen as supportive. However, their proposed ideas and methods are different from ours. For instance, rather than revealing the nature of the preference for large output norms in (conventional top-k) routing, as we analyze, CompeteSMoE Pham et al. (2024) designs a norm-based router to introduce this tendency manually; MoEfication Zhang et al. (2021) splits MLP layers of a dense model to construct MoE experts, while our study highlights that the neurons of an expert can be seen as tiny experts. Moreover, many of our observations are novel, such as the correlation between the router embedding matrix and the expert weight matrix, as well as the equivalence between parameter and output measurement for experts. Therefore, we believe that our work offers valuable insights into MoE LLMs for the community.
@@ 8 Conclusion
In this paper, we initially attempt to investigate the inner working mechanisms of MoEs by studying the parameters and outputs of four different MoE models. We summarize our empirical observations and propose practical suggestions across various aspects. While it is premature to conclude whether MoEs genuinely learn heterogeneous experts, some of our experiments indicate that specific architectural designs (e.g., the number of experts) and training frameworks may facilitate expert specialization. We hope this work can provide inspiring insights and serve as a valuable foundation for future research on MoE and other modular architectures.
@@ Appendix A Model Selection
Our experiments are conducted on Mixtral 8x7B, DeepSeekMoE, and Grok-1. We choose these models due to their widespread use and impressive performance across various domains. Additionally, these models are complementary in several crucial attributes, such as training scheme, activation functions, top-k settings, and the number of experts, as listed in Tab 1 and Tab 4. This allows for a comparative analysis with controlled variables and encompasses a wide range of parameter sizes, from rather small (16B) to relatively huge (314B). Hence, we believe that the findings derived from these four models are fairly robust, despite the limited number of models examined.
@@ Appendix B Extended Related Work
MoE LLMs. MoEs have garnered significant attention in recent years due to their ability to efficiently scale model capacity with minimal computational overhead. Most current transformer-based MoE LLMs adopt a typical architecture design that replaces the original FFN with multiple expert networks and a sparse gating network Wei et al. (2024a); Wu et al. (2024); Dai et al. (2024); Xue et al. (2024); Jiang et al. (2024); Zoph et al. (2022). JetMoE Shen et al. (2024) and ModuleFormer Shen et al. (2023) incorporate Mixture of Attention Heads Zhang et al. (2022) into their model, achieving further sparsity. A recent survey Cai et al. (2024) provides a comprehensive review of both the algorithmic and system design aspects of MoEs. For this study, we select four representative candidates among current open-sourced MoE LLMs for analysis to gain intriguing insights.
@@ C.1 Matrix-level
Figure 6: Projection of expert matrices in 2D space. Each layer contains three graphs, corresponding to $W_{\text{up}}$, $W_{\text{act}}$, and $W_{\text{down}}$, respectively. For DeepSeek, the indices of the removed outliers are listed on top of each graph.
@@ C.2 Neuron-level
Figure 7: Projection of expert neurons in 2D/3D space. Each layer contains three graphs, corresponding to $W_{\text{up}}$, $W_{\text{act}}$, and $W_{\text{down}}$, respectively.
To project the neurons into a 2D or 3D space, each row vector of $W_{\text{up}}$ and $W_{\text{act}}$, or each column vector of $W_{\text{down}}$, is treated as a single data point. Standardization is then applied, following by PCA. The visualization of the principal components is illustrated in Fig. 7. Different colors refer to neurons belonging to different experts.
Common. The vast majority of neurons gather in the low-dimensional space. In some layers, the distribution of neurons forms a special shape, such as a cross or a thick line, which appears the most often for $W_{\text{down}}$, followed by $W_{\text{up}}$, and finally $W_{\text{act}}$. Compared to ellipses, these shapes indicate that the neurons are relatively more similar.
@@ Appendix D Averaging Expert Neurons
To investigate expert correlation at the neuron level, the averaging approach simply averages the rows (for $W_{\text{up}}$ and $W_{\text{act}}$) or the columns (for $W_{\text{down}}$) of the weight matrices and then calculates the similarity of the resulting vectors across experts. Fig. 8 displays the graphs.
Common. The heat maps of $W_{\text{up}}$ and $W_{\text{down}}$ are nearly identical to those presented in § 4.1.1. Yet the similarities of $W_{\text{act}}$ significantly increase.
Mixtral and Mistral. The dark crosses sometimes disappear. In the figures for $W_{\text{act}}$, the similarities between the experts and the Mistral FFN are often lower than the similarities among the experts themselves (i.e., $S_{\text{ee}}>S_{\text{ef}}$), which is contrary to previous observations. This can happen if the expert neurons in different positions are alike. For instance, given three vectors $f=(0,0)$, $e_{1}=(1,0)$, and $e_{2}=(0,1)$, the vector similarity $S_{e_{1}e_{2}}$ is lower than $S_{e_{1}f}$ and $S_{e_{2}f}$. If averaging the elements, we have $\bar{f}=(0)$, $\bar{e}_{1}=(0.5)$, and $\bar{e}_{2}=(0.5)$, then $S_{e_{1}e_{2}}$ becomes the highest.
DeepSeek. The growth of the $W_{\text{act}}$ similarity values is directly proportional to the layer depth.
Gork. In the heat map of $W_{\text{act}}$, dark crosses frequently appear in various positions.
@@ Appendix E Gate Embedding
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Layer Mixtral Mixtral-22 DeepSeek Grok
0 0.82 -0.01 0.89
1 -0.44 0.23 0.75 -0.10
2 0.26 0.62 0.78 -0.28
3 0.54 0.76 0.71 0.66
4 0.48 0.47 0.77 0.52
5 0.70 0.49 0.77 0.37
6 0.84 0.28 0.69 0.28
7 0.74 0.69 0.73 0.17
8 0.42 0.66 0.66 0.51
9 0.66 0.84 0.66 0.84
10 0.53 0.59 0.63 0.28
11 0.32 0.61 0.60 0.30
12 0.14 0.55 0.54 0.46
13 0.51 0.48 0.60 0.14
14 0.66 0.62 0.56 0.00
15 0.40 0.59 0.58 0.54
16 0.39 0.68 0.53 0.32
17 0.53 0.65 0.55 0.30
18 0.35 0.66 0.57 0.10
19 0.17 0.72 0.57 -0.17
20 0.51 0.77 0.58 0.24
21 0.63 0.67 0.62 0.58
22 0.36 0.65 0.62 0.46
23 0.51 0.35 0.62 0.14
24 0.48 0.30 0.68 0.10
25 0.66 0.13 0.62 0.00
26 0.81 0.27 0.58 -0.10
27 0.63 0.22 0.46 -0.26
28 0.73 0.38 -0.66
29 0.75 0.52 -0.41
30 0.84 0.38 -0.83
31 0.57 0.45 -0.76
32 0.37 -0.24
33 0.28 -0.53
34 0.72 -0.46
35 0.69 0.14
36 0.34 0.17
37 0.46 -0.46
38 0.31 -0.17
39 0.26 -0.26
40 0.48 -0.70
41 0.41 0.17
42 0.46 0.00
43 0.33 -0.17
44 0.44 -0.22
45 0.50 0.14
46 0.43 -0.47
47 0.34 -0.44
48 0.50 -0.17
49 0.42 -0.14
50 0.43 0.17
51 0.51 0.22
52 0.67 0.10
53 0.32 0.33
54 0.68 -0.24
55 0.20 -0.57
@@
Table 5: Pearson correlation coefficients ($R$) of the paired dataset $(X,Y_{\text{act}})$.
Since the gate embedding matrix $W_{g}$ determines the gate decision, there may be a relationship between $W_{g}$ and the experts. To investigate this, we measure the similarities between the gate embedding vectors, $W_{g}[n,:]$ for $n\in[1,N]$, and compare them with the neuron-level averaging heat maps of experts presented in § 4.1.2. The qualitative analysis of the combined graphs shown in Fig. 8 is detailed in this section. The table containing the $R$ values for each layer (Tab. 5) is appended at the end.
Mixtral. Focusing on the heat maps of $W_{g}$, the similarities typically range from 0.2 to 0.4, with a noticeable increase in the last layer. Moreover, dark crosses are rarely found. Surprisingly, the patterns in the heat maps of $W_{g}$ and of expert neurons in $W_{\text{act}}$ are partially alike in some layers. This implies that the way a gate selects experts might be relevant to how an expert activates its neurons.
DeepSeek. Unlike the almost all-zero heat maps of $W_{\text{up}}$ and $W_{\text{down}}$, the similarities of gate neurons sometimes exceed 0.4. In addition, the heat maps of $W_{g}$ and $W_{\text{act}}$ show similar patterns. However, the overall similarities of $W_{g}$ decrease with depth while the similarities of $W_{\text{act}}$ gradually grow. This indicates that as the layer depth increases, the gate “looks” at the input feature in more diverse ways when assigning scores to different experts, even as the neuron activations of the experts become more similar.
Grok. Both dark and bright crosses commonly exist in the heat maps of $W_{g}$, whose patterns are similar to those of $W_{\text{act}}$. Specially, their patterns show opposite color tendency (i.e., deep color positions in one heat map becomes light color in another) starting form the intermediate layers. The similarities of $W_{g}$ decrease when the layer depth increases, except for the last few layers.
@@ Appendix F Additional Datasets
Figure 9: Average similarity heat maps of expert output features using (1) the long sequence, (2) WikiText-103, (3) GSM8K, and (4) Magicoder-Evol-Instruct-110K. The tick numbers refer to the expert indices. “F” and “SE” denote the Mistral FFN and the DeepSeek shared expert, respectively.
@@ Appendix H Intermediate States of Experts
While § 5.1 focused on the final outputs of experts, we continue our analysis here by examining their intermediate outputs to examine the inner states of the experts. Given an input $x$, the intermediate state of an expert refers to the output of $\sigma(W_{\text{act}}x)\in\mathbb{R}^{d_{\text{hid}}}$, where $\sigma$ denotes an activation function. These intermediate vectors control the activation of neurons, so we simply record them for analysis with the short input used. Mixtral, Mistral, and DeepSeek utilize SiLU as the activation function, while Grok adopts GeLU. Fig. 12 depicts the magnitude of the vectors for Mixtral across three tokens.
@@
Mixtral and Mistral. For a single token, we found that, on average, the absolute activation value of 99.6% elements in each expert exceeds 0.001 after applying the SiLU activation function. This high ratio indicates that the vast majority of neurons in an expert are activated. In Fig. 12, some vertical lines across all experts are commonly found, meaning that the $W_{\text{act}}$ matrices of different experts assign similar activation values to neurons with the same indices. In addition, the magnitude of the intermediate states grows along with layer depth, which aligns with the observation in § 5.1.